help with stat:

A software developer wants to re-do the survey in #1 with the goal of estimating the proportion of NKU students who carry an iPhone using a 95% confidence interval. The developer wants her estimate to be within ± 3% of the true proportion. Calculate the minimum sample size she must gather to achieve that level of accuracy. Show work for partial credit.
Answer #1

Given CI Level = 95%

p = 0.5 (as initial value of proportion is not given we assume it as 0.5)

ME = 0.03

z-value of 95% CI = 1.9600

n = (z/ME)^2*p*(1-p)

= (1.96/0.03)^2*0.5*0.5

= 1067.07

Ans: 1067